### v1

# Characterizing the Complexity of Social Robot Navigation Scenarios

Abstract
Social robot navigation algorithms are often demonstrated in overly simplified scenarios, prohibiting the extraction of practical insights about their relevance to real-world domains. Our key insight is that an understanding of the inherent complexity of a social robot navigation scenario could help characterize the limitations of existing navigation algorithms and provide actionable directions for improvement. Through an exploration of recent literature, we identify a series of factors contributing to the complexity of a scenario, disambiguating between contextual and robot-related ones. We then conduct a simulation study investigating how manipulations of contextual factors impact the performance of a variety of navigation algorithms. We find that dense and narrow environments correlate most strongly with performance drops, while the heterogeneity of agent policies and directionality of interactions have a less pronounced effect. This motivates a shift towards developing and testing algorithms under higher-complexity settings.
I Introduction
Although it is a time of rich research activity in the social navigation field [36, 15, 54], researchers often evaluate their methods with strong assumptions that ultimately do not transfer to the real world. Recent surveys on social robot navigation [36, 15, 54] have highlighted systematic problems that stem from the lack of concrete benchmarking protocols. Much of the literature considers sparse or slowly navigating human crowds, which give rise to relatively non-interactive encounters that are easy to handle by classical navigation algorithms. Hence, it remains unclear whether modern approaches are truly effective. However, it is not straightforward to define a rigorous benchmark because there is a lack of community consensus over appropriate experiment design, which would need to capture a multitude of aspects such as the spatial arrangement and roles of all agents, the briefing that human users receive before the experiment, robot design specifications, metrics evaluating robot performance and human impressions, etc.
In this paper we argue that to establish effective benchmarks for social robot navigation, the research community should better understand and control the dimensions of problem complexity. Fields like theoretical computer science, learning theory, and motion planning seek to establish clear definitions of complexity [2, 65, 13] that help define the limits of existing algorithms and motivate directions for research. However, robotics is a synthetic science [25], combining many elements from other fields, which makes it harder to formulate similarly succinct and relevant definitions for complexity. Social robot navigation [36] is no exception to this complication, as it requires the consideration of multiple disciplines, including motion planning, machine learning, control, and human factors, among others. We choose to define complexity using an empirical approach: a dimension of Complexity is a factor that causes a drop in Performance of an automated method relative to average humans (Fig. 1). Here, Performance refers to any relevant metric accepted by the community.
Figure 1: We investigate how different parameters defining a social robot navigation scenario impact the performance of recent algorithms. We show that while many algorithms successfully handle a wide range of scenarios, their performance drops at a region of Complexity that is representative of real-world situations that humans effectively handle [19].
To instantiate this analytical framework, we review the evaluation practices of recent literature and extract a set of factors contributing to the complexity of a social robot navigation scenario, distinguishing contextual factors from robot-related ones. We then conduct an extensive simulated study to understand how robot-independent, contextual factors impact algorithm performance. We find that Density and Environment Geometry have the strongest correlation with performance decline across a wide range of algorithms. We also find that the scenarios most frequently explored in prior work tend to be lower complexity, and allow simple, reactive methods to perform effectively. Our findings underscore the need for algorithms that explicitly account for diverse, high Complexity scenarios, and benchmarks that explore performance in scenarios that have high Complexity across multiple factors.
II Related Work
Recent surveys on social robot navigation [36, 15, 54, 16] emphasize the need for a standardization of benchmarking practices. Mavrogiannis et al. [36] observe the lack of rigorous statistical characterization of robot performance and human impressions, and emphasize the need for better repeatability. Francis et al. [15] collect a list of principles that a social navigation system should take into account, and propose guidelines for evaluating them. Gao and Huang [16] delve into the characterization of methodologies adopted by researchers and practitioners. Singamaneni et al. [54] provide a taxonomy of the literature, capturing aspects such as the robot type, the algorithmic approaches used, and the social navigation context. This work complements prior surveys by analyzing the hardness of the scenarios employed in the evaluation practices of published approaches. By understanding the explored territory of complexity, we aspire to provide directions for integrating more complex settings in benchmarking practices that more closely resemble what a robot would experience in the real world.
A major challenge in standardizing benchmarking in social navigation lies in the design of experiments, and specifically in balancing repeatability and emergence of natural interactions. Mavrogiannis et al. [36] classify real-world experiments into three classes: demonstrations (e.g., [57, 28, 29]), lab studies (e.g., [37, 35, 46, 42, 31]), and field studies (e.g., [59, 58, 6]), advocating the need for the extraction of statistical insights across social navigation experiments. However, the cost of conducting real-world experiments has motivated the experimentation within virtual environments [5]. Borrowing tools from graphics and the simulation of crowds [64, 20], several works have developed photorealistic simulators [21, 5, 62], while others have focused on finding quantitative metrics that capture aspects of performance previously only understood through qualitative metrics [66, 34, 53]. However, as demonstrated by Fraichard and Levesy [14], simulation is often introduced in evaluation practices with shortcomings. Common assumptions in simulated evaluations include the robot being invisible to humans, humans being simulated as homogeneous overly submissive agents, the scenarios considered being generated at random, and often in open spaces with no walls or obstacles. Such assumptions result in surprising observations, such as that a “blind” robot moving straight can outperform state-of-the-art agents [36]. Motivated by these observations, we contribute an experiment design that gives rise to more realistic crowd conditions by e.g., designing heterogeneous agents, and considering the manipulation of contextual variables like the environment geometry, the crowd directionality and density.
Another challenge in designing benchmarks is the lack of understanding of what makes a social navigation scenario hard. In prior work, the crowd density [59, 36, 55] has often been used as a proxy for complexity. However, density on its own does not capture the complexity of the motion coupling between closely interacting agents. To capture that, metrics like Path Irregularity [18] and the Topological Complexity Index [12, 34] quantify aspects of geometric and topological richness of agents’ trajectories. Our work is relevant to these approaches but it seeks to extract a more fundamental understanding of how the parameters defining a social navigation scenario impact its hardness. We are inspired by the Complexity definitions that more theoretical fields employ [2, 65, 13]. While our investigation is empirical, we see this as a first step towards the definition of a formal representation of Complexity.
III Factors of Complexity in the Literature
To approach a characterization of the Complexity of a social navigation scenario, we investigated the benchmarking practices of prior work, searching for relevant papers using the keywords “social” and “crowd” in proceedings from the ICRA, IROS, CoRL, and RSS conferences from 2015 to 2024. Additionally, we reviewed work cited in recent surveys [36, 15, 54, 16]. We organized the characteristics of evaluation methodologies and found a helpful breakdown of the factors characterizing the evaluation of each scenario could be through two categories: contextual (see Fig. 2) and robot-related factors. We tabulate the results in Table I.
TABLE I: Complexity factors in the literature. Directionality: passing (P), crossing (C), Circle Crossing (CC), Random (R); Density: Low $[0,0.1)$, Medium $[0.1,0.25]$, High $(0.25,\infty)$ $agents/m^{2}$; Environment: Small room $[0,25)$, Medium room $[25,100]$, large room $(100,\infty)$ $m^{2}$.
Paper Density Directionality Environment Robot Size $(mm^{3})$ $v_{max}(m/s)$ Kinematics Sensors Viewpoint
Angelopoulos et al. [1] Low P Hallway 425x480x1210 0.85 DD N/A Egocentric
Bennewitz et al. [4] Low N/A Office 520x520x1180 0.65 N/A Stereo, LIDAR Egocentric
Burgard et al. [6] Medium N/A Museum N/A 0.80 N/A LIDAR Egocentric
Ciou et al. [10] Low P Hallway 381x455x217 1.20 DD LIDAR, RGB-D Egocentric
Chen et al. [9] Medium P, C Hallway N/A N/A DD N/A N/A
Katyal et al. [22] Low P Med. Room 1100x500x610 1.60 Quad Motion Capture Overhead
Kim and Pineau [23] N/A P Med. Room N/A N/A N/A RGB-D Egocentric
Kirby [24] Low P Hallway 520x520x1180 0.70 N/A Laser scanner Egocentric
Liu et al. [29] Medium N/A Med. Room 354x354x420 0.50 DD LIDAR Egocentric
Liu et al. [27] Medium P, C Outdoor N/A N/A DD LIDAR Egocentric
Liu et al. [28] Low P, R Med. Room 354x354x420 0.65 DD RGB-D Egocentric
Lo et al. [31] Low P L. Room N/A 1.50 Omni Motion Capture Overhead
Matsuzaki et al. [33] Medium P, C L. Room N/A 1.20 LIDAR N/A Egocentric
Matsuzaki et al. [32] High P, C L. Room N/A 0.50 N/A N/A N/A
Mavrogiannis et al. [35] Medium P, C Sm. Room N/A 0.75 Omni Motion Capture Overhead
Mavrogiannis et al. [34] Medium P, C Med. Room 508x660x1580 0.83 Omni Motion Capture Overhead
Mun et al. [38] Low P L. Room 354x354x420 0.65 DD LIDAR Egocentric
Mustafa et al. [39] Low P Med. Room 504x430x250 2.00 DD N/A N/A
Nguyen et al. [40] Low P Hallway 810x320x640 4.80 Ackermann LIDAR Egocentric
Oh et al. [41] Low P, C L. Room 504x430x250 2.00 DD LIDAR, Stereo Egocentric
Pacchierotti et al. [42] Low P Hallway N/A 0.60 N/A LIDAR Egocentric
Paez-Granados et al. [44] High P, C Outdoor N/A N/A N/A LIDAR, RGB-D Egocentric
Peddi et al. [45] Medium P Sm. Room 960x793x296 0.50 Omni Motion Capture Overhead
Poddar et al. [46] Medium P, C Sm. Room N/A 1.50 Omni Motion Capture Overhead
Qin et al. [47] Low P, C L. Room N/A 0.85 DD LIDAR Egocentric
Qiu et al. [48] Medium P, C Hallway 354x354x420 0.65 DD LIDAR, RGB-D Egocentric
Sathyamoorthy et al. [49] High N/A Hallway 354x354x420 1.00 DD LIDAR, RGB-D Egocentric
Shiomi et al. [50] N/A N/A L. Room 600x600x1200 0.75 N/A LIDAR Overhead
Silva et al. [51] Low P Med. Room 425x480x1210 0.55 DD Motion Capture Overhead
Singamaneni et al. [52] Low P Doorway 668x668x1640 1.00 N/A LIDAR Egocentric
Tai et al. [57] Low P, C M. Room 281x306x141 0.26 DD LIDAR, RGB-D Egocentric
Thrun et al. [58] Medium N/A Museum N/A 0.70 N/A LIDAR Egocentric
Trautman et al. [59] High P, C Med. Room N/A 0.30 N/A Stereo Overhead
Truong and Ngo [60] Low P Med. Room N/A N/A N/A LIDAR, RGB-D Egocentric
Tsai and Oh [61] Medium P, C Road 990x670x390 1.00 DD LIDAR Egocentric
Wang et al. [69] Medium CC L. Room N/A N/A N/A RGB-D Egocentric
Wang et al. [68] Low CC L. Room N/A N/A N/A RGB-D Egocentric
Xie and Dames [70] Medium N/A Field 354x354x420 0.50 DD LIDAR Egocentric
Yang et al. [71] Low P, R Hallway 700x400x500 3.30 Quad Motion Capture Overhead
Yao et al. [72] Low P, C Hallway 354x354x420 0.65 DD LIDAR, RGB-D Egocentric
III-A Contextual Factors
By contextual, we refer to factors are robot agnostic, i.e., they are related to the environment the robot is deployed in.
Density. Density is often used as a proxy for the Complexity of a scenario [59, 36, 55]. We report it in the standard form of $agents/m^{2}$, although we note this does not account for the variable size of agents. Wherever a Density was not explicitly given or computable, we approximated based on supplemental footage to complete Table I.
Directionality. The directions in which humans encounter the robot contribute to the difficulty of the collision-avoidance task [67]. We identified four cases: Passing: Agents move parallel, paths do not intersect; Crossing: Agents move perpendicular, paths intersect; Random: Agent starts and goals are randomly sampled; Circle Crossing: Agent start and goals are sampled on opposite sides of the circumference of a circle.
Environment Geometry. We find that most often, evaluations take place in either hallways or medium to large office rooms, which place no constraints on the agents’ movements.
Policy Mixture. The way humans navigate next to the robot is another complicating factor. Cooperative agents assume partial responsibility for collision avoidance, which simplifies the robot’s task. In contrast, when rushing, being distracted, or changing intentions, humans may pose greater challenges to a robot [46]. We found that most real-world studies instruct participants to navigate cooperatively, and most simulation-based studies make use of cooperative crowd simulators like Social Force [20] and ORCA [63].
III-B Robot Factors
These factors refer to details of the robot hardware platform, and include the robot footprint, mass, maximum speed, and sensing capabilities, among others.
Robot Size. Most recent evaluations use robots with a smaller form factor than those in previous works, which allows other agents to maneuver around them more easily.
Maximum Speed. Consistently, the maximum speed of the robot is much lower than a natural human walking speed [46], due to hardware or safety constraints. When reviewing supplementary video material, however, we observed that this often resulted in the humans moving to avoid the robot, while the robot maintained a nearly straight trajectory.
Kinematics Most often, we see differential drive robots used in real world experiments. This is seldom matched in simulation training and testing, in which it is frequently assumed the robot is holonomic [9, 8, 36].
Sensors. Most works use limited sensing capabilities, leading to a more challenging evaluation. We note that some of the potential negative effects of limited sensing (false detections, occlusion) are more problematic in scenarios with higher numbers of pedestrians, and might not be stress-tested in many of the evaluations we covered.
Viewpoint. We differentiate between overhead, in which sensors have a bird’s eye view of the scene, mitigating issues such as occlusions, and egocentric, in which sensors are placed on the robot.
IV Investigating the Complexity of Social Robot Navigation Scenarios
We propose a definition of a Social robot navigation scenario based on parameters related to the factors of Complexity identified in Sec. III. We then describe an experiment design that investigates the performance of a variety of navigation algorithms under different scenarios.
IV-A Social Navigation Scenario
Consider a robot navigating next to $n\geq 1$ human agents in a workspace $\mathcal{W}\subseteq\mathbb{R}^{2}$ with a set of static obstacles $\mathcal{W}_{obs}\subseteq\mathcal{W}$. The robot starts from an initial configuration $s_{R}$ and moves towards a goal $g_{R}$ by following a policy $\pi_{R}$ whereas humans are navigating from their initial configurations $s_{i}$ towards their goals, $g_{i}$ by following a policy $\pi_{i}$, $i\in\mathcal{N}$; agents’ goals are unknown to one another. The robot occupies an area $A_{R}\in\mathcal{W}$, and each human occupies an area $A_{i}\in\mathcal{W}$. The objective of the robot is to reach its destination while avoiding collisions with static obstacles and abiding by social norms. We define a social navigation scenario as a tuple:
$\mathcal{S}=\left(n,A_{R},A_{i:n},s_{R},g_{R},s_{i:n},g_{i:n},\pi_{i:n},\mathcal{W}_{obs}\right)\mbox{.}$ (1)
We denote by $\pi_{i}$ the true policy for agent $i$, capturing the way they make decisions based on their objectives as well as behavioral and contextual aspects of their navigation profile.
IV-B Experiment Design
We design scenarios of varying complexity by manipulating each of the factors in Sec. III in isolation.
Scenario configurations. We first define a base scenario $\mathcal{S}_{b}$ defining with $n=15$, $A_{R},A_{i:n}=\pi(0.3)^{2}$, $s_{R}=(5,1)$, $g_{R}=(5,9)$, $\pi_{i:n}=ORCA,SFM,v_{pref}=1.0m/s$, $\mathcal{W}_{obs}=\{[0,w]\times[0,l]\}^{\mathsf{c}}=\{[0,10]\times[0,10]\}^{\mathsf{c}}$; $s_{i:n},g_{i:n}$ are sampled from passing and crossing. We then modify a single variable in each experiment, from low to high intensity:
  • Density: {0.05, 0.10, 0.15, 0.20, 0.25, 0.30, 0.35}. We choose this range following prior work [59, 36].
  • Directionality: {Passing only, crossing only, passing and crossing, circle crossing, random start/goal}. We found that prior evaluations frequently involve passing scenarios with one or two humans, although many papers set up simultaneous passing and crossing scenarios. We also find that while circle crossings are frequently used in simulation to force Complex interactions between several agents [7, 8, 36, 56], they do not appear as often in real robot evaluations.
  • Policy Mixture: {SFM only, ORCA only, Mix 1, Mix 2, Mix 3}, where Mix 1 contains 8 ORCA and 7 SFM agents, Mix 2 contains 5 ORCA, 5 SFM, 2 CV, 3 Static agents, and Mix 3 contains 4 ORCA, 4 SFM, 4 CV, 3 Static agents, respectively. We add increasing numbers of inattentive agents at higher intensity to model Complex real-world scenarios in which the robot must navigate among both cooperative and uncooperative agents simultaneously [33].
  • Environment Width: {4.5, 4.0, 3.5, 3.0, 2.5, 2.0, 1.5}. , which reflects the reported width of most hallways (e.g., [24, 71, 42]) in Table. I.
Fig. 2 illustrates the configurations considered in our study.
Refer to caption   Refer to caption   Refer to caption
a Density
Refer to caption   Refer to caption   Refer to caption
b Width
Refer to caption   Refer to caption   Refer to caption
c Directionality
Refer to caption   Refer to caption   Refer to caption
d Agent Policy Mixture
Figure 2: Left to right: Scenarios of increasing complexity for each of the complexity factors considered (d-p). In the leftmost Directionality figure green agents are crossing and blue agents are passing, while different colors represent different policies in Policy Mixture.
Algorithms. We propose to evaluate the change in complexity of a given scenario using the performance of several navigation algorithms. Specifically, we employ:
  • Relational Graph Learning (RGL) [8]: a reinforcement learning based approach that models pedestrian interactions using a graph neural network.
  • Social GAN Model Predictive Control (MPC-SGAN) [17]: An approach integrating a recent crowd motion prediction model into an MPC framework, using the approach and implementation of Poddar et al. [46].
  • Constant Velocity Model Predictive Control (MPC-CV): Identical to MPC-SGAN, but instead using CV motion predictions for each agent.
  • Social GAN Model Predictive Path Integral (MPPI-SGAN): SGAN integrated into an MPPI controller.
  • Reactive Planner (RP): A myopic planner that attempts to avoid collisions while navigating towards the goal.
  • Optimal Reciprocal Collision Avoidance (ORCA) [63]: A multiagent collision avoidance method that guarantees collision free movement among ORCA agents.
  • Social Forces Model (SFM) [20]: A physics-inspired model of crowd motion.
  • Constant Velocity (CV): An unreactive agent that moves with constant velocity toward the goal.
Metrics. Based on literature  [16, 36, 15], we use the following metrics to measure performance, collected only from Successful trials:
  • Success: The average number of trials in which the ego-agent reaches its goal without collision.
  • Time to goal: The average time to goal across trials.
  • Distance to agent: The Minimum Distance to the nearest other agent during a trial, averaged across trials.
  • Path Irregularity [18]: The amount of unnecessary turning per unit path length, measured in $\frac{rad}{m}$, calculated as $\sum_{Path}\frac{Rotation-Min.\>rotation\>needed}{Path\>length}$.
Hypotheses. We expect that scenarios of higher Complexity will pose greater navigation challenges, and this will be reflected in significant performance drops across all algorithms. Additionally, as Density is often used as a proxy for Complexity as a whole, we anticipate that it will have the strongest correlation with performance drop. Based on these expectations, We formulate the following hypotheses:
  • H1. Increasing the intensity of each of the four complexity factors (Density, Directionality, Environment Geometry, Policy Mixture) independently decreases performance with respect to (wrt) collected metrics.
  • H2. The negative correlation between Complexity and performance in H1 will be strongest for Density.
IV-C Implementation Details
We generate 500 scenarios for each condition within each experiment and report metrics as averages across all 500 trials. We fix the random seed to ensure each method experiences the same scenarios. To simulate continuous crowd motion, agents are randomly assigned new collision free starts and goals upon reaching their goal. For the MPC methods, we used the formulation of Poddar et al. [46]. For MPC-SGAN and MPPI-SGAN, we use a checkpoint pretrained on the Zara portion of the UCY dataset [26]. For the ORCA and SFM methods, we used the default parameter settings. The reactive planner uses the same action space of Chen et al. [8], and selects the collision free action that minimizes its distance to the goal in the following timestep. All methods handle static obstacle collisions by setting the component of the action in the direction of the colliding obstacle to 00. For specific details of the training procedure for RGL, and parameter settings and tuning for the MPC and MPPI methods, as well as our enhanced version of CrowdNav, see our GitHub site at https://github.com/fluentrobotics/ComplexityNav. Videos of our experiments can be found at https://youtu.be/-ir12VoSCkY.
IV-D Results
Figure 3: Performance of methods across our experiments. Rows indicate experiments and columns correspond to different evaluation metrics. Each point represents the mean over 500 experiments; shaded regions indicate standard deviation. Mix 1 has 7 SFM and 8 ORCA agents. Mix 2 has 5 SFM, 5 ORCA, 2 CV, and 3 static agents. Mix 3 has 4 SFM, 4 ORCA, 4 CV, 3 static agents.
  Refer to caption   Refer to caption
Fig. 3 summarizes our experimental results, organized by experiment types (rows) and metrics (columns). Based on this data, we investigate the validity of our hypotheses:
H1. We find a statistically significant correlation between Density and Success rate and Minimum Distance to agent ($\rho=-0.867,-0.735$, $p<0.001$ using Spearman’s r test), and a moderate correlation between Density and average Time and Path Irregularity ($\rho=0.499,0.463$, $p<0.001$). We also find a correlation between increasing width and method Success ($\rho=0.593$, $p<0.001$), and a correlation between width and Minimum Distance ($\rho=0.285$, $p<0.05$). We do not find any statistically significant correlation with average Time or Path Irregularity. Regarding Policy Mixture and Directionality, we find strong and moderate correlations respectively between higher intensity and Success ($\rho=-0.715,-0.462$, $p<0.002$), and a moderate correlation between Directionality and average Time ($\rho=-0.314$, $p<0.05$). We see no statistically significant correlation for Minimum Distance, and Path Irregularity in either case. Thus, we find that while each factor affects at least one metric, Density and Environment Geometry appear to have the strongest correlation, giving partial support for H1.
H2. We observe that the correlations between Density and other collected metrics are all stronger than those of the other factors. Thus we find support for H2.
IV-E Analysis
While Density correlates strongest with Complexity, we see that the Environment Geometry, Directionality, and agent policies all have at least some correlation with performance, and thus should be considered when assessing the Complexity of a scenario. While the support for H2 vindicates prior use of Density as a proxy for Complexity, the partial support for H1 suggests that a more accurate picture can be obtained by analyzing the rest of the contextual factors.
RGL achieves higher Success than the MPC and MPPI methods in nearly all experiments, particularly in high density scenarios. Its Success rate performance nearly mirrors ORCA (its training Policy Mixture), however with consistently higher Minimum Distances to humans. This indicates it learned similar collision avoidance behavior to ORCA, while also learning to better respect agents’ personal space.
MPC-CV, MPC-SGAN, and MPPI-SGAN performance scales poorly with moderate to high crowd density. We do observe that MPPI-SGAN consistently has higher Success rates than MPC-SGAN, showing that a stronger controller implementation does indeed lead to better performance (albeit with slightly lower Minimum Distances). The steep decrease of all three methods, however, indicates that having an inaccurate prediction model becomes increasingly problematic as the number of agent trajectories predicted increases, even with a more robust controller.
In nearly all experiments, RP and ORCA policies had low average Time, while still maintaining comparable or better Success rates to other methods. Even the CV agent was moderately Successful in lower Complexity scenarios, which matches prior experimental outcomes [36]. SFM’s high Success rates and Minimum Distances indicate that with proper tuning it could be viable as a local controller for a socially navigating robot, although more must be done to evaluate whether its effectiveness transfers out of simulation.
We do however see that the MPC methods maintain the highest Minimum Distance, while ORCA, RP, and CV have the lowest Minimum Distances. Additionally, the MPC methods, with the exception of the Environment Width experiments, maintain comparable Path Irregularity to ORCA, SFM, and RGL, showing they can maintain distance with comparably smooth paths. Thus we see that while the reactive and CV methods (with the exception of SFM) achieve efficient, collision-free navigation through other agents’ cooperation, they are worse at respecting the personal space of others compared to predictive methods.
We obtained some evidence that increased heterogeneity of scenarios is not by itself an indicator of increased Complexity. Qualitatively, we observed that passing scenarios are generally easier to navigate than crossing, and SFM agents are much more subservient than ORCA. An implicit hypothesis in our experimental ordering was that a mixture of two Directionalities or Policy Mixtures would be more Complex to navigate than either individually. We instead see improvement in the ORCA/SFM and Passing/Crossing scenarios compared to ORCA and Crossing only, which suggests the combinations are actually less Complex.
V Discussion
Conducting high-Complexity evaluations. Evaluations in the literature are typically conducted under low-Density settings (Table I). Our experiments demonstrate that most algorithms handle those well but experience significant performance reductions as Density increases. Similarly, while many works contain passing-only experiments, in our study, Passing has the highest Success rates and Minimum Distances to agents, indicating it is the easiest to navigate. Additionally, while many works consider Medium-to-Large rooms, these appear to be the simplest to navigate (Fig. 3). Finally, nearly all evaluations in recent literature focus on scenarios with exclusively cooperative agents, which our experiments showed are easiest. These practices hinder the extraction of helpful insights; for example, had we not manipulated the Complexity factors towards the upper extremes, the severe performance drops at high density experienced by MPC-CV, MPC-SGAN, and MPPI-SGAN would not have been identified. These observations suggest that the social navigation community should shift to studying scenarios of higher Complexity to move beyond the frontier. Specifically, more evaluations in more geometrically-constrained, highly-mixed and high-traffic environments, similar to many context-rich public domains should be conducted.
Handling test-time distribution. The high Success rates of CV and RP suggest that in lower-Complexity regimes, it might be sufficient to use non-reactive classical planning techniques. As the Complexity increases, all algorithms experience a steep decline; the magnitude of the decline is related to how far out of distribution an algorithm operates. Our evaluation (Fig. 3) captures the sensitivity of data-driven approaches to their training distribution. For instance, we see that RGL, and MPC/MPPI-SGAN experience a substantial performance decline as the test-time Density moves away from the training/tuning Density ($0.05,0.10$ $agents/m^{2}$). In practice, data-driven approaches will often face distribution shift at deployment, which motivates the development of algorithms for overcoming distribution shifts to be an important direction for future research [30, 73]. Alternative approaches include techniques for proactive, expressive motion generation [34, 3, 66] or techniques that embrace graceful touch [43], similar to how human crowds resolve close interactions in train stations or airports [19].
Reproducing Complexity. In our literature review, we found that the amount of details provided varied substantially among papers, leaving us to approximate values based on figures or supplementary material, when available. We believe that sharing more precise details related to the Complexity factors we identified will be an important step towards better reproducibility in social navigation.
Simulating complex settings. While realistically simulating pedestrian-robot interactions is challenging [14, 36], we view it as an essential tool for testing high-Complexity settings that are difficult to safely replicate in the real world. This requires revisiting conventional assumptions, such as that humans are non-reactive to the robot [38, 29, 68, 32, 9, 8], which is unrealistic since the robots used (see Table I) have a large enough footprint to be observed. Thus, simulation should focus on visible robot settings, leveraging metrics and considerations of users’ perceptions [34, 53, 66] to compare algorithms’ social performance.
Limitations. Our analysis is a step towards understanding the Complexity of social navigation scenarios; an important next step would be to validate our observations with real-world experiments. While our evaluation involves standard metrics [36, 15], it lacks a user-centered perspective. Future work will incorporate metrics mapping robot behavior to users’ impressions [11, 34, 66]. We tested algorithms for which we found open-source implementations but additional algorithms could be included. We strived for a fair comparison by tuning all algorithms wrt the same criteria. While tuning can always be improved, the key takeaway remains: there is a steep drop in performance as Complexity increases. Finally, our analysis only explored the impact of contextual factors on robot performance; future work will look into the robot-related factors of kinematics, sensors, and viewpoints.